The proportion of eligible voters in the election who will vote for the ANC is assumed to be 0.55 in Gauteng. What is the probability that in a random of 500 Gauteng voters less than 0.49 say they will vote for the ANC?


1
Expert's answer
2021-10-18T07:37:02-0400

Let "X" be a binomial rv based on "n" trials with success probability "p." Then if the binomial probability histogram is not too skewed, "X" has approximately a normal distribution with "\mu=np" and "\sigma^2=npq."

In practice, the approximation is adequate provided that both "np\geq 10" and "nq\geq 10."


Given "n=500, p=0.55, q=1-p=1-0.55=0.45."

Check "np=500(0.55)=275>10, nq=500(0.45)=225>10."


Then "\mu=np=275, \sigma^2=500(0.55)(0.45)=123.75."


"X\sim N(275, 123.75)"


Given "p_1=0.49"



"0.49(500)=245""P(X<0.49)=P(Z<\dfrac{245-275}{\sqrt{123.75}})""\approx P(Z<-2.6928)\approx0.0035"